# When Algorithms Meet Artists: Topic Modeling the AI-Art Debate, 2013–2025

### v1

@@ Abstract
As generative AI continues to reshape artistic production and alternate modes of human expression, artists whose livelihoods are most directly affected have raised urgent concerns about consent, transparency, and the future of creative labor [9, 1]. However, the voices of artists are often marginalized in dominant public and scholarly discourse [2]. This study presents a twelve-year analysis, from 2013 to 2025, of English-language discourse surrounding AI-generated art. It draws from 439 curated 500-word excerpts sampled from opinion articles, news reports, blogs, legal filings, and spoken-word transcripts. Through a reproducible methodology, we identify five stable thematic clusters and uncover a misalignment between artists’ perceptions and prevailing media narratives. Our findings highlight how the use of technical jargon can function as a subtle form of gatekeeping, often sidelining the very issues artists deem most urgent. Our work provides a BERTopic-based methodology [6] and a multimodal baseline for future research, alongside a clear call for deeper, transparency-driven engagement with artist perspectives in the evolving AI-creative landscape.
@@ 1 Introduction
Artists have not been passive observers of this shift. Survey research by Lovato et al. [9] reveals that many practitioners regard non‑consensual dataset scraping as a serious threat and advocate for full transparency around AI training inputs. This survey identifies four dominant frames in practitioners’ own words:
@@
Do artists perceive generative AI as a material threat to their livelihoods?
@@
Do they also recognise practical or creative benefits?
@@
Do they demand detailed disclosure of training data and model provenance?
@@
How do they assign rights among original artists, model developers, and end users?
While Lovato et al. provides a space for artists to voice concern over non-consensual scraping and call for full dataset transparency, it remains unclear to what extent public discourse foregrounds or overshadows these artist-identified concerns. Prior computational studies have primarily focused on short‑form content such as social media [8] or formal scholarly literature [17], leaving the broader, multimodal public discourse on generative AI and art largely unmapped. Our study builds on this literature by offering the first multimodal, consensus-based map of generative AI art discourse, identifying both clear topic categories and overlapping themes that appropriately interpret public conversation. Specifically, in this paper, we analyze twelve years (2013–2025) of English-language discourse across blogs, podcasts, legal documents, news articles, and academic papers. Through leveraging transformer‑based embeddings, k‑means clustering, and BERTopic [6], together with consensus‑based dimensional reduction, we aim to address the following central questions:
  • Topic structure: What coherent clusters structure public discourse on generative AI art?
  • Artist concerns: How do these clusters relate to artist-identified themes Threat, Utility, Transparency, and Ownership [9]?
  • Temporal dynamics: How have these themes shifted in prominence over time, particularly around major technical releases or legal events?
We hypothesize that public discourse partially aligns with the four artists’ frames (Threat, Utility, Transparency, Ownership) identified by Lovato et al. [9]. Specifically, technical narratives will dominate and drown out artists’ concerns of Transparency during major model releases.
Contributions
We identify five thematically distinct topics and situate them along two continuous axes (D1: Artistic to Technical Engineering; D2: General Market to Legal & Regulatory).
We quantify year-by-year shifts in cluster prevalence against major generative-AI milestones (2014–2025).
We map those empirical clusters back to artist perceptions, revealing where public discourse amplifies versus overshadows artist voices.
@@ 2.1 Data Collection and Preprocessing
We assembled a multimodal corpus of English-language discourse on AI-generated art spanning from 2013 through early 2025. Using a curated set of search terms (appendix) we gathered content from a range of sources, including academic journals, conference proceedings, policy briefs, legal rulings, blog posts, opinion pieces, podcasts, YouTube, TEDx presentations, and artist interviews. All spoken media were transcribed using automated speech recognition (ASR) and manually reviewed for accuracy. Each source was tagged by media type (e.g., article, podcast, peer-reviewed paper) and categorized by subgenre (e.g., legal commentary, artist panel, solo talk) to support metadata-aware analysis.
To prepare the data for natural language processing, we applied a standardized preprocessing pipeline. This included lowercasing, removing URLs, dates, numerals, and formatting artifacts, followed by lemmatization and stop-word removal. To minimize selection bias, we also excluded search query terms such as “AI,” “art,” “artwork,” and “technology.” Each document was segmented into non-overlapping 500-word sections, producing a final dataset of 439 text chunks of relatively uniform length for downstream embedding and analysis.
@@ 2.4 Consensus Dimensions via Procrustes-Aligned UMAP
To extract continuous dimensions of discourse, we projected each BERTopic output into two dimensions using UMAP [10] with fixed random states. We then aligned the five UMAP outputs to a common reference using Procrustes analysis [5]. This alignment process showed minimal variation across embeddings, with disparities ranging from 0.025 to 0.045, indicating strong geometric consistency. By averaging across aligned outputs, we derived two consensus axes: $D1_{consensus}$ and $D2_{consensus}$. These capture key semantic trends. The first axis ranges from community-oriented, artistic discourse to technical, engineering language. The second axis spans from conversational, general-market framing to formal legal and intellectual property concerns.
@@ 2.5 Downstream Analyses
2.6 Statistical Alignment with Artist Survey
To quantify how closely public discourse mirrors the four artist-identified frames from Lovato et al. (2024) [Threat (RQ1), Utility (RQ2), Transparency (RQ3), and Ownership (RQ4)], we transformed each section’s consensus coordinates ($D1$, $D2$) into four nonnegative “alignment weights.”
We began by linearly rescaling each axis to [1,1][-1,1]. From there, we then computed raw quadrant scores by multiplying the appropriate half-axis projections (e.g. Threat == (D1)times(D2)(-D1)\\ times(-D2)), and normalized those so that for each section the four weights sum to one. This yields per-section alignment scores: $p_{threat}$, $p_{utility}$, $p_{transparency}$, $p_{ownership}$.
@@ 3.2 Topics over Time
Figure 2: Proportion of each topic by year (2013–2025), with milestone annotations (light gray).
@@ 3.3 Readability Gradient
3.4 Statistical Alignment with Artist Survey
@@
3.5 Alignment with Artist Frames
@@ 3.6 Linguistic Metrics
Taken together, these oscillations in diversity map onto the technology and legal milestones we’ve already traced in Figure 2. Peaks in diversity often align with new narrative needs—e.g., fresh metaphors for threat (2016) and community responses (2021)—whereas more “settled” storylines around market sizing or legal norms show less lexical variety.
@@
Figure 5 presents a Pearson-correlation heatmap over our section-level features (sentiment, subjectivity, readability, lexical diversity, log-article-word-count, $D1_{consensus}$, $D2_{consensus}$, and prob_consensus). A few key patterns emerge:
  • Lexical Diversity $\leftrightarrow$ Flesch–Kincaid Grade (r=0.61r=-0.61): As FK grade goes up, diversity goes down. In practice, the highest-grade passages (post-graduate readability) live in highly specialized subfields—either deep engineering or formal legal discourse—which tend to recycle a smaller core of technical or legal terms rather than drawing on a broad vocabulary.
@@
  • $D1_{consensus}\leftrightarrow$ Lexical Diversity (r=0.26r=-0.26): Moving toward the technical end of $D1$ correlates with lower diversity, echoing our readability finding that high-$D1$ sections demand higher grade-levels to parse (they use a narrow band of jargon repeatedly).
  • $D2_{consensus}\leftrightarrow$ Lexical Diversity ($r=+0.40$): Moving up the legal/IP end of $D2$ correlates with higher diversity—legal and policy debates often invoke a broader palette of statutory, philosophical, and economic terms.
@@
  • $D1_{consensus}\leftrightarrow$ Sentiment (r=0.15r=-0.15): Technical passages skew slightly more negative—perhaps reflecting critical or cautionary assessments of model limitations.
  • $D2_{consensus}\leftrightarrow$ Sentiment ($r=+0.22$) and (r=0.36r=-0.36) with polarity: Legal/regulatory sections trend less positive and less subjective overall, consistent with the formal, prescriptive tone of policy and copyright discourse.
@@
The lack of strong correlations between prob_consensus and any other metric ($|r|<0.10$) indicates that our consensus-based topic assignments are not simply a byproduct of section length, readability, or sentiment—they represent a distinct semantic signal.

### v2

@@ Abstract
As generative AI continues to reshape artistic production and alternate modes of human expression, artists whose livelihoods are most directly affected have raised urgent concerns about consent, transparency, and the future of creative labor [9, 1]. However, the voices of artists are often marginalized in dominant public and scholarly discourse [2]. This study presents a twelve-year analysis, from 2013 to 2025, of English-language discourse surrounding AI-generated art. It draws from 439 curated 500-word excerpts sampled from opinion articles, news reports, blogs, legal filings, and spoken-word transcripts. Through a reproducible methodology and three routes of evidence, including a distributional test, in-text frame geometry, and machine learning predictions, we identify five stable thematic clusters and triangulate on a misalignment between artists’ perceptions and prevailing media narratives. Our findings highlight how the use of technical jargon can function as a subtle form of gatekeeping, often sidelining the very issues artists deem most urgent. Our work provides a BERTopic-based methodology [6] and a multimodal baseline for future research, alongside a clear call for deeper, transparency-driven engagement with artist perspectives in the evolving AI-creative landscape.
@@ 1 Introduction
Artists have not been passive observers of this shift. Survey research by Lovato et al. [9] reveals that many practitioners regard non‑consensual dataset scraping as a serious threat and advocate for full transparency around AI training inputs. Building on the practitioner’s survey responses, we have developed four dominant frames that synthesize the concerns raised:
@@
Do most artists see Generative AI models as a threat to art workers?
@@
Do most artists see Generative AI art models as either a positive or neutral development in the field of art?
@@
Do most artists believe that creators of AI models should disclose the art and images used to train their models in detail?
@@
Do most artists believe that the work of Generative AI models and its derivatives should be considered the property of the artist whose style is recognizable and mimicked by the Generative AI art model and not the property of the creators of AI models?
While Lovato et al. (2024) provide a space for artists to voice concerns about non-consensual scraping and to call for full dataset transparency, it remains unclear to what extent broader public discourse foregrounds or overshadows these artist-identified priorities. Prior computational studies focus largely on short-form social media [8] or formal scholarly literature [17], leaving the multimodal public sphere around AI and art comparatively unmapped. We address this gap by constructing a multimodal, consensus-based map of generative-AI art discourse spanning blogs, podcasts, legal filings, news, and academic articles (2013–2025).
Methodologically, we leverage transformer-based embeddings, k-means clustering, and BERTopic [6], together with consensus-based dimensional reduction, to situate each section of text in a two-dimensional semantic space (D1: Artistic Practice $\rightarrow$ Technical Engineering; D2: General Market $\rightarrow$ Legal & Regulatory). We use this space to examine three central questions:
  1. (i)
    Topic structure: What coherent clusters organize public discourse on generative AI art?
  2. (ii)
    Artist concerns: How do those clusters relate to the artist-identified frames - Threat, Utility, Transparency, Ownership - reported by Lovato and et al. [9]?
  3. (iii)
    Temporal dynamics: How do these clusters and frames wax and wane around major technical releases and legal milestones?
To support and validate the consensus topic structure and the D1/D2 axes, we adopt a triangulated design with three independent routes of evidence that jointly address these questions:
Route 1 - Distributional alignment.
We map each section’s D1/D2 coordinates to frame weights (Threat, Utility, Transparency, Ownership) and test whether the corpus-level frame proportions align with Lovato’s survey proportions. This route targets artist concerns directly and, in aggregate, informs topic structure.
Route 2 - Sentence-level frame inference.
Using survey-faithful prototype phrases, we compute prototype similarity per section and track frame intensities over time. This route adds temporal dynamics and provides an in-text check that the axes carry the intended frame geometry.
Route 3 - Geometric predictive validity.
We test whether section locations on the D1/D2 plane are predictable from language (n-grams plus readability/lexical-diversity features) via ensemble classifiers. Successful prediction supports the interpretability and stability of the axes and topics (topic structure) and helps explain artist-concern variation by language and style.
Contributions.
  • We identify five interpretable discourse topics and place them on two consensus axes that capture stylistic and semantic spectra salient to AI–art debates.
  • We quantify year-by-year shifts in these topics relative to major technical and legal milestones (2014–2025).
  • We link public frames to artist frames via three convergent routes of evidence, revealing where public discourse amplifies versus over- or under-represents the priorities voiced by artists.
  • We release a reproducible pipeline for consensus topic modeling and axis validation that generalizes to longitudinal, multimodal corpora.
Hypothesis.
We hypothesize partial alignment between public discourse and artists’ framing: specifically, technical and utilitarian narratives will dominate coverage, especially around major model releases, while Transparency (and, at times, Ownership) will be under-represented relative to artists’ stated priorities.
@@ 2.1 Data Collection and Preprocessing
We assembled a multimodal corpus of English-language discourse on AI-generated art spanning 2013 through early 2025. Using Google Search and a curated query list (Appendix Table S1), we targeted high-visibility, public-facing sources, excluding social-media platforms and forums, and systematically sampled items that appeared on the first results page for each query. To reduce personalization effects, searches were performed in a logged-out browser with a cleared cache; the date of access was recorded for each item. When Google’s “AI Overview” panel appeared, we followed its citations to the underlying sources and included only those original documents; the AI-generated summary text itself was not used as corpus content.
The resulting corpus covers a broad range of genres: peer-reviewed journal and conference articles, policy briefs and legal rulings, news and opinion pieces, blog posts, and long-form spoken content (podcasts, YouTube/TED/TEDx talks, artist panels and interviews). Spoken items were transcribed using automated speech recognition (ASR) and manually reviewed for basic quality control (e.g., speaker names, technical terms). Each source was coded with metadata for media type (e.g., article, podcast, peer-reviewed paper) and subgenre (e.g., legal commentary, artist panel, solo talk) to support downstream, metadata-aware analyses.
Preprocessing followed a standardized NLP pipeline tailored to mixed-genre text. We lower-cased, removed URLs, formatting artifacts, and standalone numerals/dates, then applied lemmatization and stop-word removal. To mitigate query-driven selection bias, we also filtered a small domain-specific stoplist comprising the most generic search terms (e.g., AI, art, artwork, technology) used to retrieve the corpus. Documents were segmented into non-overlapping 500-word sections to standardize unit length across media types, yielding 439 analytic sections for embedding and modeling. This length normalization improved comparability across short opinion pieces, long essays, and transcripts.
We performed basic deduplication (removing exact URL duplicates and mirrored copies) and retained the canonical link where possible. All sources were publicly available at the time of collection; quotations reported in the article adhere to fair-use limits.
We note two sampling constraints that we treat as part of the phenomena under study: (i) first-page bias in web search privileges high-salience narratives and established outlets; and (ii) search results can vary by time and geography. Our triangulated design and transparent query list are intended to make these choices reproducible while acknowledging their implications for coverage.
@@ 2.4 Consensus Dimensions via Procrustes-Aligned UMAP
To extract continuous dimensions of discourse, we projected each BERTopic output into two dimensions using UMAP [10] with fixed random states. We then aligned the five UMAP outputs to a common reference using Procrustes analysis [5]. This alignment process showed minimal variation across embeddings, with disparities ranging from 0.025 to 0.045, indicating strong geometric consistency. By averaging across aligned outputs, we derived two consensus axes: $D1$ and $D2$. These capture key semantic trends. The first axis ranges from community-oriented, artistic discourse to technical, engineering language. The second axis spans from conversational, general-market framing to formal legal and intellectual property concerns.
@@ 2.5 Downstream Analyses
2.6 Triangulation Inference: Three Routes of Evidence
Frame Alignment
To quantify how closely public discourse mirrors the four artist-identified frames from Lovato et al. (2024) [Threat (RQ1), Utility (RQ2), Transparency (RQ3), and Ownership (RQ4)], we triangulated on our results and conclusions using three routes of evidence. This triangulation intentionally mixes distributional, temporal, and predictive evidence. If all three converge, we gain stronger confidence about alignment and misalignment than any single method alone.
Route 1: Distributional Test
We transformed each section’s consensus coordinates (D1, D2) into four nonnegative ’alignment weights’ by linearly rescaling each axis to [1,1][-1,1]. From there, we then computed raw quadrant scores by multiplying the appropriate half-axis projections (e.g. Threat == $(-D1)\times(-D2)$), and normalized those so that for each section the four weights sum to one. This yields per-section alignment scores: $p_{threat}$, $p_{utility}$, $p_{transparency}$, $p_{ownership}$.
@@ 2.6 Triangulation Inference: Three Routes of Evidence
Route 2: Sentence-level Frame Inference
We generated three paraphrased sets of frame phrases and five prototype sentences per frame to infer per-section frame intensities. We used the four RQs (outlined in Introduction) as the prototypes per frame and computed similarity with cosine similarity. The similarities of each frame were analyzed across the years of the corpus. We could then measure any major changes in the ratio of similarity across the four frames and if changes correlated with any milestone events.
Route 3: Geometric Prediction Models
We train a series of random-forest classifiers on the top 12,000 n-grams (one to three) along with a series of linguistic features - including lexical diversity, readability score (FK grade), subjectivity score, and others - to predict geometric placement on our D1 and D2 axes. We use a series of different numbers of trees (from 500 to 5,000) per classifier separately on three different target variables (y). The targets included: quadrant location on the D1 axis, quadrant location on the D2 axis, and quadrant location on the 2-dimensional D1/D2 space. For each round of classifiers, we ensembled the predictions across the classifiers and then found average feature importance for the top 10 features across all rounds.
@@ 3.2 Topics over Time
Figure 2: Proportion of each topic by year (2013–2025), with milestone annotations with milestones in light gray.
@@ 3.3 Readability Gradient
3.4 Triangulation of Inference: Three Routes of Evidence
@@
3.4.1 Route 1: Distributional Comparison
@@
Figure 3: Distribution of prototype similarities over years 2013-2025 with key milestones in light gray.
Route 2: Prototype Similarity
Figure 4 plots the mean cosine similarity between each section summary and the prototype phrase per frame. Utility remains highest ($\sim 0.50-0.60$), Threat and Ownership are mid‑range ($\sim 0.30-0.40$), and Transparency is near zero across the entire period, indicating chronic under‑emphasis. Wilcoxon rank‑sum tests (Table 6) reveal a post‑2024 increase in Ownership (p=0.043) and decrease in Threat (p=0.0035); other year cuts are not significant. Combined with Route-1, these results confirm that artists’ central demand, transparency, remains systematically muted in public discourse.
Route 3: Geometric Prediction
Random‑forest ensembles predict D2 quartiles at 79% weighted-accuracy (78% macro-accuracy) and 2‑D quadrants at 82% weighted-accuracy (81% macro-accuracy), with D1 quartiles at 62% weighted-accuracy (62% macro-accuracy). Top features mix content n‑grams (e.g. neural, image) and style metrics (log FK grade, lexical diversity, subjectivity), confirming that register is part of what defines the axes. In particular, readability and lexical diversity reliably separate legal/technical texts from market/community coverage, echoing Fig. 4’s accessibility gradient (below). Across all three rounds of classifiers, readability and lexical diversity were always in the top 10 most important features and usually within the top six (out of >>12,000 features).
@@ 3.5 Linguistic Metrics
Taken together, these oscillations in diversity map onto the technology and legal milestones we’ve already traced in Figure 2. Peaks in diversity often align with new narrative needs–e.g., fresh metaphors for threat (2016) and community responses (2021)–whereas more “settled” storylines around market sizing or legal norms show less lexical variety.
@@
Figure 5 presents a Pearson-correlation heatmap over our section-level features (sentiment, subjectivity, readability, lexical diversity, log-article-word-count, $D1$, $D2$, and prob_consensus). A few key patterns emerge:
  • Lexical Diversity $\leftrightarrow$ Flesch–Kincaid Grade (r=0.61r=-0.61): As FK grade goes up, diversity goes down. In practice, the highest-grade passages (post-graduate readability) live in highly specialized subfields–either deep engineering or formal legal discourse–which tend to recycle a smaller core of technical or legal terms rather than drawing on a broad vocabulary.
@@
  • $D1\leftrightarrow$ Lexical Diversity (r=0.26r=-0.26): Moving toward the technical end of $D1$ correlates with lower diversity, echoing our readability finding that high-$D1$ sections demand higher grade-levels to parse (they use a narrow band of jargon repeatedly).
  • $D2\leftrightarrow$ Lexical Diversity ($r=+0.40$): Moving up the legal/IP end of $D2$ correlates with higher diversity—legal and policy debates often invoke a broader palette of statutory, philosophical, and economic terms.
@@
  • $D1\leftrightarrow$ Sentiment (r=0.15r=-0.15): Technical passages skew slightly more negative—perhaps reflecting critical or cautionary assessments of model limitations.
  • $D2\leftrightarrow$ Sentiment ($r=+0.22$) and (r=0.36r=-0.36) with polarity: Legal/regulatory sections trend less positive and less subjective overall, consistent with the formal, prescriptive tone of policy and copyright discourse.
@@
The lack of strong correlations between the probability of consensus and any other metric ($|r|<0.10$) indicates that our consensus-based topic assignments are not simply a byproduct of section length, readability, or sentiment—they represent a distinct semantic signal.
@@ Appendix A Appendix: Additional Tables & Figures
Table 6: Wilcoxon rank-sum pre/post milestone tests for prototype frames. Two-sided $z$ statistics and raw $p$ values; $q$ values are Benjamini–Hochberg FDR within frame (12 tests). Direction encodes change in the post period relative to pre. Significance: ** $q<0.05$; * $p<0.05$ (unadjusted); $\dagger$ $p<0.10$.
(a) Transparency
Frame Year $z$ Milestone $p$ $q$ Direction Sig
Transparency 2014-2025 NA >>0.05 NA NA
(b) Threat
Frame Year $z$ Milestone $p$ $q$ Direction Sig
Threat 2014-2020 NA - >>0.1 NA NA
Threat 2021 -1.95 2021 5.072e-02 0.152 $\downarrow$ post $\dagger$
Threat 2022 -1.93 StableDiffusion4 5.377e-02 0.152 $\downarrow$ post $\dagger$
Threat 2023 -1.46 Andersen v. Stability AI 1.455e-01 0.225 $\downarrow$ post
Threat 2024 -2.92 GPT-4o 3.514e-03 0.042 $\downarrow$ post **
Threat 2025 -1.89 EU AI Act5/1st AI art auction 5.931e-02 0.152 $\downarrow$ post $\dagger$
(c) Ownership
Frame Year $z$ Milestone $p$ $q$ Direction Sig
Ownership 2014-2020 NA >>0.1 NA NA
Ownership 2021 1.66 2021 9.712e-02 0.325 $\uparrow$ post $\dagger$
Ownership 2022 1.62 StableDiffusion4 1.062e-01 0.325 $\uparrow$ post
Ownership 2023 1.32 Andersen v. Stability AI 1.875e-01 0.325 $\uparrow$ post
Ownership 2024 2.02 GPT-4o 4.316e-02 0.325 $\uparrow$ post *
Ownership 2025 1.31 EU AI Act5/1st AI art auction 1.895e-01 0.325 $\uparrow$ post
(d) Utility
Frame Year $z$ Milestone $p$ $q$ Direction Sig
Utility 2014-2017 NA >>0.1 NA NA
Utility 2018 -1.85 Christie’s/BERT/GPT-1 6.409e-02 0.256 $\downarrow$ post $\dagger$
Utility 2019 -1.95 GPT-2 5.103e-02 0.256 $\downarrow$ post $\dagger$
Utility 2020 -1.45 GPT-3 API 1.477e-01 0.443 $\downarrow$ post
Utility 2021 -1.90 2021 5.682e-02 0.256 $\downarrow$ post $\dagger$
Utility 2022-2025 NA >>0.1 NA NA
Table 7: Route 3 results: confusion matrices (counts and row-normalized) and averaged feature importances for the RF ensembles.
(a) D1 quadrants (D1_q): confusion matrices
Pred 1 Pred 2 Pred 3 Pred 4
True 1 16 2 3 1
True 2 4 13 3 1
True 3 4 4 9 4
True 4 1 0 6 15
Pred 1 Pred 2 Pred 3 Pred 4
True 1 0.727000 0.091000 0.136000 0.045000
True 2 0.190000 0.619000 0.143000 0.048000
True 3 0.190000 0.190000 0.429000 0.190000
True 4 0.045000 0.000000 0.273000 0.682000
(b) D2 quadrants (D2_q): confusion matrices
Pred 1 Pred 2 Pred 3 Pred 4
True 1 22 0 0 0
True 2 3 14 1 3
True 3 1 3 13 4
True 4 0 1 3 18
Pred 1 Pred 2 Pred 3 Pred 4
True 1 1.000000 0.000000 0.000000 0.000000
True 2 0.143000 0.667000 0.048000 0.143000
True 3 0.048000 0.143000 0.619000 0.190000
True 4 0.000000 0.045000 0.136000 0.818000
(c) 2D quadrant labels (quad4_sign): confusion matrices
Pred 1 Pred 2 Pred 3 Pred 4
True 1 20 1 1 0
True 2 7 12 1 1
True 3 2 0 14 5
True 4 2 0 1 19
Pred 1 Pred 2 Pred 3 Pred 4
True 1 0.909000 0.045000 0.045000 0.000000
True 2 0.333000 0.571000 0.048000 0.048000
True 3 0.095000 0.000000 0.667000 0.238000
True 4 0.091000 0.000000 0.045000 0.864000
Table 8: Top predictive features across Route-3 ensembles. Importances are min-max normalized within ensemble. Results are from splitting the D1 and D2 axes into 4 quadrants as targets (Y), one at a time. Quad-4 results are from using the 2-dimensional quadrants as targets (Y).
feature D1_q D2_q Quad-4 Mean importance
think 0.380 0.918 1.000 0.766
like 0.481 1.000 0.807 0.762
know 0.725 0.708 0.791 0.742
thing 0.663 0.576 0.945 0.728
log_fk_grade 0.636 0.509 0.615 0.587
lexical_diversity 0.675 0.480 0.463 0.539
image 0.822 0.213 0.560 0.532
tb_subjectivity 0.862 0.379 0.322 0.521
not 0.521 0.480 0.444 0.482
prob_consensus 0.722 0.349 0.335 0.469
yeah 0.323 0.670 0.387 0.460
go 0.381 0.544 0.426 0.451
tb_polarity 0.400 0.457 0.409 0.422
tw_sent 0.635 0.274 0.301 0.403
kind 0.365 0.410 0.399 0.391
Table 8: Top predictive features across Route-3 ensembles. Importances are min-max normalized within ensemble. Results are from splitting the D1 and D2 axes into 4 quadrants as targets (Y), one at a time. Quad-4 results are from using the 2-dimensional quadrants as targets (Y).